OpenStudy (anonymous):

Find all real values of x such that f(x) =0. f(x)=x^3-x^2-4x+4

OpenStudy (anonymous):

x(x^2-4)-(x^2-4) = (x^2-4)(x-1)

OpenStudy (anonymous):

(x^2-4)(x-1) ==> x=2 or x=-2 or x=1

OpenStudy (anonymous):